# svg_figures/

### v1

@@ 5.1 Future directions
  • Broader Model Classes: Model collapse has been extensively studied in image generation domains. It remains an open question whether similar loss-based strategies can be adapted to diffusion models [19] and generative adversarial networks (GANs) [31], where diversity collapse is also a central concern.
@@
  • Collapse-Aware Generation: While our approach targets the training process, model collapse can also be addressed during the generation phase. Combining our training-side solution with decoding-aware techniques such as [15] may yield complementary benefits, improving the quality and diversity of generated data further.
@@ Appendix A Focal Loss
Focal Loss (FL), originally introduced for addressing class imbalance in image segmentation tasks [32], emphasizes hard, misclassified examples. It is defined as:
@@
Recent work [33] has shown FL’s effectiveness in language modeling, outperforming CE on commonsense reasoning and math tasks. In the context of this work, FL aligns with our objective of mitigating model collapse by emphasizing rare token learning and preserving the tail of the output distribution.
@@ E.1 Synthetic datasets
4 commercials 26, 1993, Jordan participated in a three @-@ point contest at Madison I’ve been gambling since I can remember. That’s how I made wrote a book claiming he had won \$ 1 @.@ 25 million from Jordan on sports bets. Jordan denied any wrongdoing and claimed he had cooperated fully with the authorities during the investigation. On March 31, 1993, He said, ” I have never gambled in my life. I don ’t know why anyone would think that I would lie about that. Obviously I wouldn ’t such a thing.” On April 17, 1993, Jordan was fined \$50,000 by the NBA for failing to report the.
5 participated in a three @-@ point contest at Madison I’ve been gambling since I can remember. That’s how I made wrote a book claiming he had won \$1 billion.@ 25 million from Jordan on sports bets. Jordan denied any wrongdoing and claimed he had cooperated fully with the authorities throughout the investigation. On March 31, 1993, He said, ” I have never gambled in my life. I don ’t know why anyone would think that I would lie about that. Obviously I wouldn’t such a thing.” On April 17, 1993, Jordan was fined \$50,000 by the NBA. Where did the confusion on the 50 mil fine to Jordan come from: probably: The Bulls were fined \$ 30 @,@ 000 for the game : \$,@ 000 for failing to report the impromptu number change to the NBA and \$ 5 @,@ 000 for Jordan wearing different shoes.
Appendix F GMM experiment algorithm

### v3

@@ 5.1 Future directions
  • Broader Model Classes: Model collapse has been extensively studied in image generation domains. It remains an open question whether similar loss-based strategies can be adapted to diffusion models (Ho et al., 2020) and generative adversarial networks (GANs) (Goodfellow et al., 2014), where diversity collapse is also a central concern.
@@
  • Collapse-Aware Generation: While our approach targets the training process, model collapse can also be addressed during the generation phase. Combining our training-side solution with decoding-aware techniques such as Drayson et al. (2025) may yield complementary benefits, improving the quality and diversity of generated data further.
Appendix A Appendix
@@ Appendix A Focal Loss
Focal Loss (FL), originally introduced for addressing class imbalance in image segmentation tasks Lin et al. (2018), emphasizes hard, misclassified examples. It is defined as:
@@
Recent work Rege Cambrin et al. (2024) has shown FL’s effectiveness in language modeling, outperforming CE on commonsense reasoning and math tasks. In the context of this work, FL aligns with our objective of mitigating model collapse by emphasizing rare token learning and preserving the tail of the output distribution.
@@ E.1 Synthetic datasets
4 commercials 26, 1993, Jordan participated in a three @-@ point contest at Madison I’ve been gambling since I can remember. That’s how I made wrote a book claiming he had won \$ 1 @.@ 25 million from Jordan on sports bets. Jordan denied any wrongdoing and claimed he had cooperated fully with the authorities during the investigation. On March 31, 1993, He said, " I have never gambled in my life. I don ’t know why anyone would think that I would lie about that. Obviously I wouldn ’t such a thing." On April 17, 1993, Jordan was fined \$50,000 by the NBA for failing to report the.
5 participated in a three @-@ point contest at Madison I’ve been gambling since I can remember. That’s how I made wrote a book claiming he had won \$1 billion.@ 25 million from Jordan on sports bets. Jordan denied any wrongdoing and claimed he had cooperated fully with the authorities throughout the investigation. On March 31, 1993, He said, " I have never gambled in my life. I don ’t know why anyone would think that I would lie about that. Obviously I wouldn’t such a thing." On April 17, 1993, Jordan was fined \$50,000 by the NBA. Where did the confusion on the 50 mil fine to Jordan come from: probably: The Bulls were fined \$ 30 @,@ 000 for the game : \$,@ 000 for failing to report the impromptu number change to the NBA and \$ 5 @,@ 000 for Jordan wearing different shoes.
Appendix F Gaussian Mixture Models (GMM)
Unlike deep learning models discussed earlier, Gaussian Mixture Models (GMMs) are not based on neural networks. Instead, they are trained by maximizing the likelihood of the input data under a mixture of $N$ Gaussian components, parameterized by means and variances. GMMs are central to previous theoretical and empirical studies of model collapse Shumailov et al. (2024a), and extending our method to this setting underscores its broad applicability.
Our intuition mirrors that of the LLM experiments: we aim to prevent the model from overfitting to high-probability modes while preserving low-probability, diverse samples. To achieve this, we simply exclude data points that fall within the top $\alpha$-percentile of likelihoods under the current model. In our experiments (Figure 12), we set $\gamma=0.8$.
To simulate recursive data generation in a non-neural setting, we adapt our filtering strategy to Gaussian Mixture Models. Algorithm 2 describes the procedure: starting from an initial dataset, a GMM is repeatedly fit to the current data and used to sample new points. At each iteration, we compute the likelihood of each sampled point under the fitted model and discard the highest-probability samples, retaining only those below the $\gamma$-percentile threshold. This ensures that subsequent datasets remain diverse and are not dominated by a small set of high-likelihood modes, thereby mirroring the mitigation strategy applied in our language model experiments.
Appendix G Variational Auto Encoders (VAE)
To demonstrate the applicability of confidence-aware loss functions beyond language modeling, we extend our analysis to the image generation domain using Variational Auto Encoders (VAEs), following the setup in Shumailov et al. (2024a). We train a VAE on the MNIST dataset Deng (2012) using the standard evidence lower bound (ELBO) objective, modified with our truncation function: